Sam invested $3m in a portfolio made of 10 different securities with portfolio’s Beta of 3.5.

One month later, Sam sold security A which Beta =1.8 for $490,000 (purchased at a cost of 400,000) and used the profit to add security C with beta =2.4.

Compute the new portfolio Beta.


1
Expert's answer
2022-01-23T15:37:34-0500

"\frac{3-0.490-0.500}{3}\times3.5+1.8\times\frac{0.490}{3}+2.5\times\frac{0.500}{3}=3.039"